OpenStudy (anonymous):

Choose the correct simplification of the expression 9/h-3

OpenStudy (zehanz):

If it is \[\frac{ 9 }{ h }-3\]you can write 3 as 3h/h:\[\frac{ 9 }{ h }-\frac{ 3h }{ h }=...\]So you can write it as one fraction.

OpenStudy (anonymous):

no its

OpenStudy (zehanz):

In that case, because there is this rule: \[a^{-b}=\frac{ 1 }{ a^b }\]you get:\[9 \cdot h^3=9h^3\]

OpenStudy (zehanz):

@thatgirl666: do you understand?

OpenStudy (anonymous):

yepp :)